# Chaos Engineering: A Multi-Vocal Literature Review

### v1

@@ 2.5.2. Synthesis and Analysis
We applied thematic analysis to synthesize insights across our research questions, ensuring that coding remained objective and independent of the data extraction structure or biases (Stewart, 2024). Full papers were subsequently imported into ATLAS.ti, a qualitative data analysis tool that allowed for direct text coding. This process adhered to established thematic analysis steps (Braun and Clarke, 2006), providing consistency and depth in our approach. We used a step-by-step approach to systematically identify codes, clusters, and classifications related to our research questions (Kumara et al., 2021):
@@
  • Pilot Study: We randomly selected twenty sources for initial analysis, with the first two authors of the paper independently analyzing each to establish preliminary codes using structural and descriptive coding approaches (Saldaña, 2021). This phase involved creating codes by conceptualizing information related to the research questions. Through iterative review and refinement, researchers adjusted codes to clarify emerging theories and form thematic clusters.
  • Inter-Rater Consensus Analysis for the Pilot Study: After finalizing the coding process, the first two authors of the paper conducted an inter-rater review, assessing each other’s codes to establish consensus on naming, types, and categorization. This collaborative effort helped to refine codes and maintain their consistency.
  • Complete Dataset Coding: The remaining sources were subsequently coded based on the agreement reached during the pilot study. The first author of this paper conducted this step, and the second author reviewed the resulting codes.
  • Inter-Rater Consensus Analysis for the Complete Study: After completing the individual coding, we performed a final inter-rater assessment to ensure the coding aligned with the established framework. Any discrepancies were discussed and resolved.
  • Generating the report: The results of this thematic analysis are presented in Sections 3,  4,  5,  6 and  7.
@@ 2.6. Replication Package
For the validation and reproducibility of this work, we have provided an extensive replication package online [^1] https://drive.google.com/drive/folders/1wOu11lB9wsxK7cQKzPL63GAjD9UiiF0s?usp=sharing. This package includes the complete compilation of sources, search queries, and thematic analysis performed using Atlas.ti [^2] https://atlasti.com/ (such as codes, groupings, quotations, and analyses).
@@ 3. Chaos Engineering: Definitions, Functionaries, and Elements (RQ1)
This section first examines practitioners’ definitions of chaos engineering. Next, it describes the phases, functional requirements, and non-functional requirements of chaos engineering. Finally, it identifies the main functional components of chaos engineering.
@@ 3.1. Definitions
Chaos engineering is a growing practice, but its definition remains inconsistent. Hence, we aim to establish a clear and shared definition by analyzing those proposed by different sources. Originally developed at Netflix, chaos engineering focuses on intentionally introducing faults within a system to assess and enhance its robustness  (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Torkura et al., 2021). However, confusion often arises when distinguishing chaos engineering from traditional testing practices. While conventional testing remains vital, chaos engineering should be seen as a complement rather than a replacement (Green, 2023). Traditional testing and monitoring are crucial, yet they may not be sufficient to expose all possible failure modes in complex systems. Chaos engineering mitigates this issue by proactively identifying the weaknesses of a system that might go unnoticed during standard testing procedures but may surface as failures in production environments (Paul, 2019; Green, 2023).
Table 6. Selected Definitions of Chaos Engineering
Source Definition
Netflix
"the discipline of experimenting on a system to build confidence in its ability to withstand turbulent
conditions in production" (Team, 2024).
Gremlin
"a disciplined approach to identifying failures before they become outages by intentionally
’breaking things’ to build more resilient systems" (Inc., 2023).
AWS
"the process of stressing an application in testing or production environments by creating
disruptive events, such as server outages or API throttling, observing how the system responds,
and implementing improvements" (Starr et al., 2022; Kahil and Arenhage, 2021).
Google "the practice of intentionally introducing failures into production systems under load" (Google Cloud, 2024).
Microsoft "intentionally injecting faults into applications in a controlled manner to simulate real-world issues" (Microsoft, 2023).
IBM
"intentionally causing failures in production or pre-production environments
to understand their impact and improve defense strategies" (IBM, 2024).
Datadog
"a discipline that encourages experimenting in production and injecting controlled failures
into the system to understand how the system will react and improve reliability" (Pulido et al., 2021).
Harness "the science of injecting failure to gauge system resiliency" (Schillerstrom, 2022).
Twilio "intentionally injecting failures to test a system’s ability to self-heal" (Wong, 2017).
Uber "the practice of injecting controlled failures directly into production environments to test resilience" (Jones, 2020).
Splunk
"the practice of assessing the resilience of production systems by intentionally introducing
failures and running experiments" (Wickramasinghe, 2023).
Figure 4. Word Cloud of Key Terms in Chaos Engineering Definitions
Table 6 shows the excerpts from the definitions provided by a subset of the reviewed literature. We also created a word cloud to identify the key terms used in these definitions, as shown in Figure  4. The key themes emphasized in these definitions include system resilience, failure injection, testing in production environments, and controlled experiments. These recurring elements highlight chaos engineering’s focus on deliberately introducing failures in a controlled manner to assess how systems behave under stress and recover from disruptions. Based on these insights, we propose the following unified definition.
@@
In line with the definition, chaos engineering relies on several guiding principles to design and carry out experiments, ensuring that systems are rigorously tested under controlled conditions (Basiri et al., 2016; Aisosa, 2023). Table 7 summarizes these principles.
Table 7. Principles and Measurement Metrics
Principle Description Articles Referenced
Establish Steady State Understand the system’s normal behavior and hypothesize how it will handle chaos (real-world disruptive incidents/failures). (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Wickramasinghe, 2023; Cotroneo et al., 2022; Nikolaidis et al., 2021; Frank et al., 2021b; Pierce et al., 2021; Frank et al., 2023; Konstantinou et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Bocetta, 2019; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Sbiai et al., 2023; Green, 2023; Krivas and Portela, 2020; Kostic, 2024; Abdul, 2024; Katirtzis, 2022; Le, 2022; Colyer, 2019; Patel, 2022)
@@
Conduct Chaos Experiments Design and execute experiments to simulate real-world failures. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Pierce et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Jernberg et al., 2020; Jakkaraju, 2020; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Patel, 2022; Abdul, 2024; Roa, 2022)
@@
Expand to Production Run experiments in production to build confidence in the system’s ability to endure adverse real-world events. (Nombela, 2023; Krishnamurthy, 2021; Moon, 2022; Simonsson et al., 2021; Pierce et al., 2021; Konstantinou et al., 2021; Park et al., 2019; Fawcett, 2020)
3.2. Functionalities of Chaos Engineering
Based on our review of academic and grey literature, we identified five key features that define an effective chaos engineering solution: experimentation hub, collaborative workflow management, observability suite, security and governance, and integration and automation (see Figure  5). The rest of this section discusses each in detail.
3.2.1. Experimentation Hub
Experimentation is a crucial aspect of a chaos engineering platform, allowing users to meticulously plan and seamlessly conduct chaos experiments (Katirtzis, 2022; Abdul, 2024; Krivas and Portela, 2020). In this phase, users first define the scope, targets, and parameters of chaos experiments with precision, ensuring a clear focus for testing (Kostic, 2024; Bremmers, 2021; Kadikar, 2023; Wu, 2021; Cotroneo et al., 2022; Katirtzis, 2022; Torkura et al., 2020). The next step is to establish the steady state of the system - baseline performance metrics and normal behavior - which serves as a reference point for evaluating the impact of disruptions (Frank et al., 2023; Konstantinou et al., 2021; Jernberg et al., 2020; Siwach et al., 2022). Based on this steady state, users form hypotheses about how the system will behave under stress, guiding both the design of experiments and the setting of resilience expectations  (Kostic, 2024; Zhang et al., 2019; Wickramasinghe, 2023; Fogli et al., 2023; Dedousis et al., 2023). Once the experiment is defined, the chaos engineering platform injects faults, such as network latency, packet loss, or server failures. This enables users to assess the weaknesses and operational readiness of a system by evaluating its fault handling capabilities (Kostic, 2024; Chen et al., 2024; Fogli et al., 2023; Zhang et al., 2023). As these faults are injected and system behavior is observed, careful planning and execution help chaos experiments effectively identify and address system vulnerabilities, ultimately strengthening overall resilience (Nikolaidis et al., 2021; Colyer, 2019).
3.2.2. Collaborative Workflow Management
Chaos engineering is a collaborative practice that brings together stakeholders such as developers, testers, operations teams, and security experts. These individuals must work closely to design, execute, and analyze experiments to identify system vulnerabilities. Through effective collaborative workflow management, teams can maintain clear communication, coordinate tasks efficiently, and ensure that all parties are informed throughout the experimentation process (Siwach et al., 2022; Dua, 2024; Green, 2023).
Facilitating Team Communication and Coordination.
Chaos engineering relies on continuous communication between relevant parties to ensure alignment with the objectives of the experiment and the expected results (Gill, 2021). By forming cross-functional working groups, participants can discuss scenarios, hypothesize outcomes, and maintain a shared understanding of the objectives of the experiment. This approach ensures that everyone stays coordinated and clear on their roles during the execution and analysis phases. For example, in a simulated database failure, developers can focus on diagnosing and resolving the issue while the operations group works to reroute traffic and maintain service availability. At the same time, security experts assess the potential vulnerabilities that the failure exposes, and testers verify the system response to failures (Patel, 2022). This collaborative process helps avoid information silos and supports collective learning, leading to better preparation for future incidents (Green, 2023; Kadikar, 2023).
Notify Stakeholders about Experiment Progress.
Keeping stakeholders informed about the status of the chaos engineering experiment is crucial to transparency and effective decision making. First, well-designed notification systems trigger alerts based on anomalies, such as unexpected error rates or system performance changes  (Bremmers, 2021). Then, these alerts are automatically sent to stakeholders through integrated notification features. These alerts enable quick issue communication and prompt responses to performance degradation or system instability. In addition, they provide real-time insight into system behaviors, facilitating adjustments during the experiment. Finally, this continuous feedback loop refines the monitoring systems and improves incident preparedness through improved alert accuracy (Shah, 2021).
Maintain Documentation and Knowledge Repositories.
Building a consistent knowledge base for future reference is essential for the long-term success and continuous improvement of chaos engineering practices (Vinisky, 2024; Kadikar, 2023). By maintaining detailed records of experiment scenarios, findings, and recommendations, organizations create a valuable foundation for troubleshooting and informed decision-making (Kadikar, 2023; Gill, 2022). This process should include standardized templates for experiments and comprehensive postmortem reports, ensuring that the documentation is consistent and thorough (Siwach et al., 2022). A well-organized knowledge base preserves essential information and fosters a deeper understanding of system behavior under stress  (Gill, 2021). These insights from previous experiments can directly improve system design and operational practices (Vinisky, 2024). Through consistent documentation and structured knowledge sharing, chaos engineering platforms ensure that lessons learned are preserved, actively shaping future experiments, and strengthening overall system performance (Torkura et al., 2020).
3.2.3. Observability Suite
Chaos engineering platforms use an observability suite to monitor and analyze systems during experiments, allowing teams to detect issues early, identify root causes, and take corrective actions (Green, 2023; National Australia Bank, 2020; Chen et al., 2022). Key features discussed in the following subsections include real-time dashboards and experiment impact analysis to identify faults’ root causes.
Provide Real-Time Dashboards.
Real-time observability depends on dashboards, metrics, traces, and logs for the visualization of experiments. (National Australia Bank, 2020; Gatev, 2021; Pulcinelli et al., 2023). First, dashboards, integrated with monitoring tools, can show critical metrics such as latency, throughput, error rates, and resource usage, allowing engineers to monitor system performance and quickly detect problems (Wickramasinghe, 2023; Manna, 2021; Fawcett, 2020; Green, 2023). Compared to performance metrics, traces can capture the flow of requests through the system, revealing interactions among components and identifying bottlenecks (Chen et al., 2024; Zhang et al., 2021a). Distributed tracing tools allow engineers to analyze component interactions and inefficiencies during chaos experiments (Konstantinou et al., 2021; Frank et al., 2021b). Finally, logs record system events, supporting diagnostic efforts, especially in security contexts (Klinaku et al., 2022; Chen et al., 2022; Green, 2023). Log analysis tools enable engineers to trace anomalies and maintain a complete audit trail  (Torkura et al., 2021; Soldani and Brogi, 2021; Vu et al., 2022).
Analyze Impact and Identify Root Causes of Faults.
After gathering and visualizing data through real-time dashboards, the next crucial step is to analyze the results of chaos experiments and identify the root causes of any faults, using each pillar of observability in the process  (Yu et al., 2021; Gianchandani et al., 2022) The metrics collected during chaos experiments are compared with baseline performance to detect deviations that may indicate vulnerabilities. These deviations help identify how well the system handled stress and where failures occurred (Ji et al., 2023). Logs and trace data provide deeper insight into how individual components interacted and responded to stress. This detailed data helps identify the source of failures and component interactions that contribute to performance problems (Green, 2023). These metrics, logs, and traces can be combined to facilitate root cause analysis and offer qualitative and quantitative insights. For example, this helps engineers determine whether failures were caused by bottlenecks, resource mismanagement, or faulty component interactions (Yu et al., 2021). By understanding the root causes of failures, teams can implement targeted improvements to enhance and prevent future failures (Gill, 2022; Nombela, 2023; Gianchandani et al., 2022).
3.2.4. Security and Governance
Chaos engineering platforms require a robust security and governance framework to ensure that experiments are conducted safely and comply with regulatory standards (Fogli et al., 2023; Kostic, 2024). This section outlines vital mechanisms that protect system integrity while adhering to industry regulations, focusing on implementing safety controls, automated rollbacks, and ensuring regulatory compliance.
Implement Safety Controls and Automated Rollbacks.
Our literature identifies several vital mechanisms, including abort conditions, automated rollbacks, and blast radius control, which ensure that experiments are conducted safely (Patel, 2022).
  • Abort conditions: These ensure that chaos experiments are immediately stopped if predefined thresholds are exceeded (Abdul, 2024; Sbiai et al., 2023). For example, the experiment is automatically terminated if performance metrics, such as latency or availability, degrade beyond acceptable levels. This prevents further damage and maintains system stability by stopping experiments before they can cause widespread problems (Torkura et al., 2020).
  • Automated rollbacks: These allow systems to revert to a known stable state after an experiment (Fawcett, 2020; Miles, 2019b). In cases where a simulated failure causes service disruption, automated rollback mechanisms quickly restore the system to its pre-experiment configuration. This reduces downtime and prevents prolonged instability in the system (Gill, 2022).
  • Blast radius control: This mechanism limits the scope of chaos experiments to isolated parts of the system (Wickramasinghe, 2023; Manna, 2021). For example, teams might start by targeting a single microservice instead of testing the entire infrastructure. Gradually expanding the blast radius allows controlled testing, ensuring vulnerabilities can be identified without risking widespread system failures (Abdul, 2024). Additionally, Gremlin  (Inc., 2023) recommends using the Rumsfeld matrix, a framework to categorize system behaviors during chaos experiments. This approach enables teams to understand better and anticipate various types of system responses by classifying them as follows:

### v2

@@ 3.5.2. Synthesis and Analysis
We applied thematic analysis to synthesize insights across our research questions, ensuring that coding remained objective and independent of the data extraction structure or biases (Stewart, 2024). The full papers were subsequently imported into ATLAS.ti  [^1] https://atlasti.com/, a qualitative data analysis tool that allowed for direct text coding. This process adhered to established thematic analysis steps (Braun and Clarke, 2006), providing consistency and depth in our approach. We used a step-by-step approach to systematically identify codes, clusters, and classifications related to our research questions (Kumara et al., 2021):
@@
  • Pilot Study: We randomly selected twenty sources for initial analysis, with the first author applying structural and descriptive coding (Saldaña, 2021) to conceptualize data relevant to the research questions (Kumara et al., 2021). The process began by extracting quotations—direct statements from the primary studies—and identifying keywords that captured their core meaning. For example (see Figure 3), from quotations referring to the need to introduce failure scenarios and monitor experiment duration, the keywords Introduce failure scenarios and Experiment duration were extracted. These were then grouped to form the code Experiment Execution. Similarly, the keywords Monitoring sidecar and Observability for diagnosis were extracted and abstracted into the code Monitoring and Observability. These codes were clustered into the broader theme Functional Requirements. This phase yielded 536 initial codes and an initial set of 28 candidate themes, which served as the foundation for subsequent refinement.
  • Inter-Rater Consensus Analysis for the Pilot Study: After completing the pilot coding, the second author reviewed the full set of codes independently, assessing the naming, scope, and categorization of each code. The discrepancies between authors were discussed and resolved collaboratively. This consensus process refined the initial 536 codes to a validated set of 421 codes, improving the clarity and consistency of the thematic structure.
  • Complete Dataset Coding: The remaining sources were subsequently coded using the validated codebook. The first author conducted the full-dataset coding, and the second author reviewed the outputs. Discrepancies identified during this phase were resolved collaboratively.
  • Inter-Rater Consensus Analysis for the Complete Study: After the complete coding process, a final inter-rater assessment was conducted to ensure consistent application of the codes across the dataset. The discrepancies in interpretation were discussed and resolved. During this phase, the initial 28 themes were critically reviewed for internal coherence and conceptual distinctiveness. Overlapped or weakly supported themes were consolidated or removed, resulting in a final set of 15 distinct, non-overlapping themes. At this stage, each finalized theme was explicitly mapped to the specific research question(s) it addressed (e.g., RQ1.1 – Functional Requirements, RQ3.2 – Best Practices), ensuring that the thematic framework remained purposeful and traceable throughout the analysis.
  • Generating the report: The results of this thematic analysis are presented in Sections 4,  5,  6,  7 and  8.
Figure 3. A representative output from our thematic analysis process, showing how quotations from academic and grey literature were distilled into keywords, abstracted into codes, and grouped under broader themes. ( Refer to caption: grey literature; Refer to caption: academic literature )
[…]
Illustration of qualitative coding and thematic abstraction from quotes to themes.
@@ 3.6. Replication Package
For the validation and reproducibility of this work, we have provided an extensive replication package online  [^2] https://doi.org/10.5281/zenodo.15696868. This includes the complete compilation of sources, search queries, and thematic analysis performed using Atlas.ti (such as codes, groupings, quotations, and analyses), as well as the associated code tables.
@@ 4. Chaos Engineering: Definitions, Functionaries, and Elements (RQ1)
This section first reviews how chaos engineering is defined across academic and industry sources. Next, it outlines its core activities and commonly associated functionalities (RQ1.1). Finally, it introduces the essential components (RQ1.2) and quality requirements (RQ1.3) that form the foundation for implementing effective chaos engineering platforms.
@@ 4.1. Definitions
Chaos engineering has garnered significant attention (Monroe, 2019) in both academia and industry, yet its definition remains fragmented and inconsistent. To address this, we analyzed a set of definitions identified during our thematic analysis process. These include representative definitions from each, summarized in Tables 7 and 8. Across industry and academic sources, several key elements consistently emerge: system resilience, failure injection, controlled experiments, and testing in production or production-like environments. These recurring elements highlight chaos engineering’s focus on deliberately introducing failures in a controlled manner to assess how systems behave under stress and recover from disruptions.
Table 7. Practitioners’ Definitions of Chaos Engineering
Definition Articles Referenced
"The practice of deliberately breaking systems in controlled ways to identify potential failure points before they occur naturally." (Tomka, 2024; Palani and Gupta, 2023b; Abdul, 2024; Green, 2023; Krivas and Portela, 2020; Vinisky, 2024)
"The methodology of creating disruptive events and stressing applications to observe system responses and implement targeted improvements." (Gill, 2021; Bhaskar, 2022; Bairyev, 2023; Kadikar, 2023)
"The technique of gauging system resiliency by intentionally causing failures under load to test self-healing mechanisms." (Patel, 2022; Gill, 2022; Bairyev, 2023; Kadikar, 2023; Kostic, 2024)
Table 8. Academic Definitions of Chaos Engineering
Definition Articles Referenced
"A disciplined practice of experimenting on a software or distributed system to build confidence in its ability to withstand turbulent or unexpected conditions in production." (Fogli et al., 2023; Torkura et al., 2020; Pierce et al., 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Jernberg et al., 2020; Naqvi et al., 2022; Vinisky, 2024; Basiri et al., 2016; Sousa et al., 2018; Alvaro et al., 2016)
"An empirical and scientific approach for learning how software systems behave under stress by conducting controlled experiments to evaluate and improve their resilience and fault-handling mechanisms." (Zhang et al., 2019; Sondhi et al., 2021; Ikeuchi et al., 2020; Malik et al., 2023)
"A resilience-testing approach involving the intentional injection of failures or disruptions into a system—especially in production or production-like environments—to uncover weaknesses and improve fault tolerance." (Poltronieri et al., 2022; Ji et al., 2023; Kesim et al., 2020; Konstantinou et al., 2021)
We followed the synthesis method proposed by Gong and Ribiere (Gong and Ribiere, 2021), drawing on conceptual evaluation criteria from Wacker (Wacker, 2004) and Suddaby (Suddaby, 2010) (see our online appendix Section 3.6 for details). A total of 48 definitions were analyzed, 27 from academic sources and 29 from grey literature. Based on this evaluation, 34 definitions were retained and 22 were excluded due to issues such as vagueness, conceptual overlap, or lack of definitional clarity. These inconsistencies highlight the difficulty of establishing a unified understanding of chaos engineering—a challenge observed in previous research, where fragmented definitions can hinder conceptual clarity and slow down the development of standardized practices (Gong and Ribiere, 2021; Degefa et al., 2021; Diaz-De-Arcaya et al., 2024; Awasthi et al., 2025). As demonstrated in those contexts, establishing a clear and shared definition can help eliminate conceptual ambiguity, enable the development of sound theory and valid constructs, and strengthen the connection between research and practice. Based on these insights, we propose the following unified definition:
@@
In line with the definition, chaos engineering relies on a set of core activities to guide the design and execution of experiments under controlled conditions (Basiri et al., 2016; Aisosa, 2023). Table 9 summarizes these activities.
Table 9. Core Activities in the Chaos Engineering Lifecycle
Activity Description Articles Referenced
Establish Steady State Understand the system’s normal behavior and hypothesize how it will handle chaos (real-world disruptive incidents/failures). (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Wickramasinghe, 2023; Cotroneo et al., 2022; Nikolaidis et al., 2021; Frank et al., 2021b; Pierce et al., 2021; Frank et al., 2023; Konstantinou et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Bocetta, 2019; Palani and Gupta, 2023a; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Sbiai et al., 2023; Green, 2023; Krivas and Portela, 2020; Kostic, 2024; Abdul, 2024; Katirtzis, 2022; Le, 2022; Colyer, 2019; Patel, 2022)
@@
Conduct Chaos Experiments Design and execute experiments to simulate real-world failures. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Pierce et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Jernberg et al., 2020; Jakkaraju, 2020; Palani and Gupta, 2023a; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Patel, 2022; Abdul, 2024; Roa, 2022)
@@
Expand to Production Run experiments in production to build confidence in the system’s ability to endure adverse real-world events. (Nombela, 2023; Krishnamurthy, 2021; Moon, 2022; Simonsson et al., 2021; Pierce et al., 2021; Konstantinou et al., 2021; Park et al., 2019; Fawcett, 2020; Willard and Hutson, 2024; Tucker et al., 2018; Basiri et al., 2016; Alvaro et al., 2016)
4.2. Functionalities of Chaos Engineering (RQ1.1)
Chaos engineering comprises a structured set of core functionalities designed to reveal system weaknesses under real-world failure conditions. Based on our thematic analysis of academic and grey literature (see Section 3.5), we identified five key functionalities that characterize how chaos engineering is typically conducted: experiment design and planning, fault execution and control, monitoring and observability, post-experiment analysis, and automation and continuous integration (see Figure 4). For full transparency, the detailed mapping of raw codes to these functionalities and supporting themes is provided in the online appendix (see Section 3.6). The remainder of this section describes each functionality, emphasizing its purpose and role in validating system resilience.
Figure 4. Key Functionalities Provided by Chaos Engineering.
[…]
Functional Requirement
4.2.1. Experiment Design and Planning
This phase establishes the conceptual and procedural foundation for chaos experimentation. It ensures that testing is deliberate, measurable, and risk-aware (Fogli et al., 2023; Simonsson et al., 2021; Patel, 2022; Vu et al., 2022). It involves three essential activities: identifying test targets and failure scenarios, defining steady-state behavior, and formulating testable hypotheses.
Identify Test Targets and Failure Scenarios
The first step is to identify components whose failure could significantly disrupt operations, such as microservices, APIs, databases, or message queues (Patel, 2022; Naqvi et al., 2022; Vu et al., 2022; Bremmers, 2021). Failure scenarios are then defined based on system architecture analysis (e.g., service dependencies, communication patterns), operational risks, or past incident reports (Fogli et al., 2023; Chen et al., 2024; Malik et al., 2023; Nikolaidis et al., 2023). Typical disruptions include network latency, packet loss, service crashes, or CPU throttling (Simonsson et al., 2021; Chen et al., 2024). These scenarios are prioritized according to their likelihood and potential business impact (Fogli et al., 2023; Naqvi et al., 2022; Siwach et al., 2022; Torkura et al., 2020).
Define Steady-State Behavior
Steady-state refers to the normal, healthy performance of the system. Teams define this using measurable indicators such as latency, error rates, throughput, or availability (Frank et al., 2023; Malik et al., 2023; Simonsson et al., 2021). These metrics act as baselines against which system behavior is compared during and after failure injection (Naqvi et al., 2022; Siwach et al., 2022; Gill, 2021). Establishing these baselines enables reliable detection of performance degradation and validation of resilience mechanisms (Naqvi et al., 2022; Inc., 2023).
Formulate Testable Hypotheses
Hypotheses describe the expected system response to injected faults. For instance, a hypothesis might state: “If service A fails, traffic will reroute to service B within 200ms without user impact” (Dedousis et al., 2023; Simonsson et al., 2021). These statements guide the design of experiments and help interpret the results. To manage uncertainty and uncover hidden risks, teams can use structured frameworks such as the Rumsfeld Matrix, which categorizes system knowledge as follows (Inc., 2023; Bocetta, 2019; Dedousis et al., 2023; Pierce et al., 2021; Malik et al., 2023; Palani and Gupta, 2023b).